Calculate the viscosity of air at (a) 273 K, (b) 298 K, (c) 1000 K. Take
σ ≈ 0.40 nm2
. (The experimental values are 173 µP at 273 K, 182 µP at 20°C,
and 394 µP at 600°C
1
Expert's answer
2020-10-26T14:48:07-0400

M = 29*10-3 kg/mol

σ = 0.4*10-18 m2 (collision cross-section)

"\eta=(M*R*T)" 0.5/("\pi"0.5*NA* σ)

(a) "\eta=(29*10" -3*8.314*273)0.5/(3.140.5*6.02*1023* 0.4*10-18) = 1.90*10-5 kg/(m*s) = 190 µP

(b) "\eta=(29*10" -3*8.314*298)0.5/(3.140.5*6.02*1023* 0.4*10-18) = 1.99*10-5 kg/(m*s) = 199 µP

(c) "\eta=(29*10" -3*8.314*1000)0.5/(3.140.5*6.02*1023* 0.4*10-18) = 3.64*10-5 kg/(m*s) = 364 µP